# MaxSup: Overcoming Representation Collapse in Label Smoothing

### v3

@@ 4.4 Visualization via Class Activation Maps
To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [Selvaraju_2019], which highlights regions most influential for each prediction.
@@ 5 Conclusion
Limitations. Prior work [muller2019does] notes that LS-trained teachers may degrade knowledge distillation [hinton2015distilling, Hu2021CVPR], and guo2024cross suggests LS accelerates convergence via improved conditioning. Examining MaxSup’s potential role in distillation and its overall impact on training dynamics would clarify these underlying effects. Recent studies [sukenik2024neural, garrod2024persistence] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
@@ Acknowledgments
This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, and the U.S. DOT PacTrans sub-center seed funding program. We also thank the anonymous reviewers for their constructive feedback and insightful suggestions, which helped improve the quality and clarity of this paper.
@@ Appendix H Extended Evaluation of Linear Transferability on Different Datasets
Table 12: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization. Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup. Following [kornblith2021better], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$. Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@ Appendix I Extended Comparison to More Label Smoothing Alternatives
We have included Confidence Penalty [pereyra2017regularizing] and Adaptive Label Smooothing with Self-Knowledge [lee2022adaptive] for an extended comparison in Table˜13. We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [gao2020towards, wei2022mitigatingneuralnetworkoverconfidence]. Specifically, the method proposed in [wei2022mitigatingneuralnetworkoverconfidence] aims to mitigate overconfidence via logit normalization during training. With its default hyperparameter settings, it achieves an accuracy of 74.32% on ImageNet with a ResNet-50, which is significantly lower than the 76.91% achieved by standard Label Smoothing. This performance aligns with that of Logit Penalty, which similarly minimizes the global $l_{2}$-norm of logits and can struggle to match baseline LS performance. We also note that these norm-based methods[gao2020towards, wei2022mitigatingneuralnetworkoverconfidence, dauphin2021deconstructing] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
@@
To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of muller2019does. Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space. We then project each sample’s penultimate-layer activation vector onto this plane. To ensure the visual clarity of the resulting plots, we randomly sample 80 images from the training or validation set for each of the three classes.
@@
We conducted an ablation study on the $\alpha$ schedule, as shown in Table˜15. The consistently high accuracy across settings demonstrates the robustness of MaxSup. The adaptive $\alpha$ schedule, adopted from [lee2022adaptive], further highlights the method’s integrity and compatibility with principled design choices.

### v4

@@ 4.4 Visualization via Class Activation Maps
To better understand how MaxSup fundamentally differs from Label Smoothing (LS) in guiding model decisions, we employ Gradient-weighted Class Activation Mapping (Grad-CAM) [29], which highlights regions most influential for each prediction.
@@ 5 Conclusion
Limitations. Prior work [23] notes that LS-trained teachers may degrade knowledge distillation [13, 14], and Guo et al. [8] suggests LS accelerates convergence via improved conditioning. Examining MaxSup’s potential role in distillation and its overall impact on training dynamics would clarify these underlying effects. Recent studies [33, 7] also show that $\ell_{2}$ regularization biases final-layer features toward low-rank solutions, raising interesting questions about whether MaxSup behaves similarly.
@@ Acknowledgments
This work was supported by the University of Washington Faculty Startup Fund, the Carwein–Andrews Fellowship, the UW GSFEI Top Scholar Award, the U.S. DOT PacTrans sub-center seed funding program, the DFG Research Unit 5336 - Learning to Sense (L2S), and the ELSA – European Lighthouse on Secure and Safe AI funded by the European Union under grant agreement No. 101070617. Views and opinions expressed are however those of the authors only and do not necessarily reflect those of the European Union or European Commission. Neither the European Union nor the European Commission can be held responsible for them.
We thank the anonymous reviewers for their helpful comments.
@@ Appendix H Extended Evaluation of Linear Transferability on Different Datasets
Table 12: Validation performance of different methods, evaluated using multinomial logistic regression with l2 regularization. Although Label Smoothing and OLS improve ImageNet accuracy, they substantially degrade transfer accuracy compared to MaxSup. Following [15], we selected from 45 logarithmically spaced values between $10^{-6}$ and $10^{5}$. Notably, the search range is larger than the search range used in Table 3, thus leading to higher overall accuracies on CIFAR10.
@@ Appendix I Extended Comparison to More Label Smoothing Alternatives
We have included Confidence Penalty [25] and Adaptive Label Smooothing with Self-Knowledge [20] for an extended comparison in Table˜13. We follow their recommended hyperparameter settings: weight coefficient is set to 0.1 for confidence penalty, and the checkpoint with the highest validation accuracy checkpoint is treated as teacher for Adaptive Label Smoothing.
To further address the novelty of our work, we additionally compared our method to recent approaches that also identify and aim to fix the issues of Label Smoothing on misclassified samples [6, 38]. Specifically, the method proposed in [38] aims to mitigate overconfidence via logit normalization during training. With its default hyperparameter settings, it achieves an accuracy of 74.32% on ImageNet with a ResNet-50, which is significantly lower than the 76.91% achieved by standard Label Smoothing. This performance aligns with that of Logit Penalty, which similarly minimizes the global $l_{2}$-norm of logits and can struggle to match baseline LS performance. We also note that these norm-based methods[6, 38, 4] are often highly sensitive to hyperparameter choices, which can limit their practical applicability.
@@
To illustrate the differences between Max Suppression Regularization and Label Smoothing, we follow the projection technique of Müller et al. [23]. Specifically, we select three semantically related classes and construct an orthonormal basis for the plane intersecting their class templates in feature space. We then project each sample’s penultimate-layer activation vector onto this plane. To ensure the visual clarity of the resulting plots, we randomly sample 80 images from the training or validation set for each of the three classes.
@@
We conducted an ablation study on the $\alpha$ schedule, as shown in Table˜15. The consistently high accuracy across settings demonstrates the robustness of MaxSup. The adaptive $\alpha$ schedule, adopted from [20], further highlights the method’s integrity and compatibility with principled design choices.